Week of 190128

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Vincent (Security), Gavin (Computing), Maarten (WLCG), Borja (Monitoring), Roberto (Storage), Renato(LHCb)
  • remote: Darren (RAL), Xavier (KIT), Marcelo (CNAF), Onno (NL-T1), Stephan (CMS)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Smooth running at 300-325k job slots
    • Some monitoring problems (missing data in old and new dashboard), data recovered
    • EOSATLAS backend migration foreseen on 12th of March: couldEOS team please declare the downtime in ITSSB so that we can circulate the URL amongst the shifters ?

  • CMS reports ( raw view) -
    • No major grid computing issue(s)
    • SAM3 and SSB stuck Wednesday to Saturday; thanks to Nikolay for clearing corrupted data/restoring updating of the dashboards!
    • Running smoothly at about 200k cores
      • 75% production, 25% analysis

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Data reconstruction for 2018 data on going
    • User jobs running and MC jobs at "full steam"

Sites / Services round table:

  • ASGC: ASGC is going to have Lunar New Year break next week (4/2~8/2). Will keep best effort operation then.
  • BNL: NC
  • CNAF: NTR
  • EGI: NC
  • FNAL: NTR
  • IN2P3: NC
  • JINR: NTR
  • KISTI: NC
  • KIT: NTR
  • NDGF: NTR
  • NL-T1: Downtime planned in 9th week - 2 days long with possible extension. More information to be provided in coming meetings.
  • NRC-KI: NC
  • OSG: NC
  • PIC: NC
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services:
    • Previously announced VOMS intervention (to remove lcg2-voms.cern.ch) cancelled .
    • OTG:0046088: LSF public deco, Wednesday 30th January 2019. (Dedicated shares handled separately.)
    • OTG:0047300: Draining HTCondor public now for CC7 upgrades. Schedule:
      • OTG:0048002 ce511, ce512, ce513, ce514 will be changed to default to CC7 at 10.00 on Tuesday 29th January (i.e. tomorrow).
      • end January 2019: 30% public/grid will be CC7 - on target.
      • end March 2019: 50% public/grid will be CC7
      • 2nd April 2019: lxplus.cern.ch alias change to CC7 (lxplus6 service will remain accessible on lxplus6.cern.ch), Default HTCondor target change to CC7 for local submission.
      • early June 2019: remainder of capacity will have been migrated.
      • (Dedicated shares handled separately.)
  • CERN storage services:
    • LHC EOS instances' backend migration to QuarkDB:
  • CERN databases: NTR
  • GGUS:
    • A new release is planned for Wed this week
      • Release notes
      • A downtime has been scheduled for 07:00-09:30 UTC
      • Test alarms will be submitted as usual
  • Monitoring:
    • Re-computation period for December availability extended since the draft reports were not notified correctly to the sites
    • Issues with the SAM3/SSB dashboards, some sites availability couldn't be computed properly, should be solved now
  • MW Officer:
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r19 < r18 < r17 < r16 < r15 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r19 - 2019-01-28 - MiroslavPotocky
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback